OpenStudy (anonymous):

Find the perimeter of a rectangular object which has a length of √75 feet and a width of √48 feet

OpenStudy (anonymous):

P=2L=2W

OpenStudy (anonymous):

L=\[\sqrt{75}\] W=\[\sqrt{48}\]

OpenStudy (anonymous):

\[P=2\sqrt{75}+2\sqrt{48}\]

OpenStudy (anonymous):

now to clean this up you can simplify more

OpenStudy (anonymous):

\[2\sqrt{25*3}+2\sqrt{16*3}\]

OpenStudy (anonymous):

\[2\sqrt{5*5*3}+2\sqrt{4*4*3}\]

OpenStudy (anonymous):

you can pull out 5 and a 4

OpenStudy (anonymous):

\[2(5)\sqrt{3}+2(4)\sqrt{3}\]

OpenStudy (anonymous):

\[10\sqrt{3}+8\sqrt{3}=18\sqrt{3}\]

OpenStudy (anonymous):

\[P=18\sqrt{3}\]